# Description length of canonical and microcanonical models

### v3

@@ Abstract
The (non-)equivalence of canonical and microcanonical ensembles is a central concept in statistical physics. Non-equivalence has recently been established in models with an extensive number of constraints, common, e.g., in network science, via a non-vanishing difference in relative entropy, corresponding to higher microcanonical log-likelihood per node. However, from a model selection perspective, comparing canonical and microcanonical models requires consideration of both log-likelihood and complexity. To compare both terms under the Minimum Description Length (MDL) principle, we compute the Normalized Maximum Likelihood (NML) of binary canonical and microcanonical models, finding that (i) microcanonical models, though higher in likelihood, are always more complex, making the choice of model non-trivial. (ii) The optimal model choice depends on the empirical values of the constraints: the canonical model performs best when its fit to observed data exceeds its uniform average fit across all data. (iii) Notably, in the thermodynamic limit the difference in description length per node vanishes for the equivalent models considered but persists otherwise, showing that non-equivalence implies extensive differences between large canonical and microcanonical models. Finally, we compare the NML approach to Bayesian methods, showing that (iv) the choice of priors, while practically uninfluential in equivalent models, becomes crucial when an extensive number of constraints is enforced, possibly leading to very different outcomes.
@@ I Introduction
Non-equivalence between canonical and microcanonical models can be formally defined in several (under mild conditions, equivalent) ways [6]. A particularly convenient definition is the measure-level one, which is rooted in information theory and involves the Kullback-Leibler (KL) divergence between the microcanonical and canonical probability distributions. According to this definition, (non-)equivalence is characterized by a (non-)vanishing relative entropy density as the system size approaches infinity. Interestingly, since the KL divergence coincides with the difference between the canonical and the microcanonical log-likelihood functions, measure-level (non-)equivalence can be inspected by simply considering the asymptotic behavior of such difference [7]. Non-equivalence has been recently shown to arise in models that involve an extensive number of constraints (e.g.,the Configuration Model for networks [7, 8, 9]) and may have practical implications for the expected macroscopic properties of the system under study (see [10] for an example concerning bipartite networks).
Whenever the two approaches are equivalent, they identify the same set of typical configurations. In these cases, preferring one formulation over the other becomes a matter of convenience (e.g.,computational efficiency). Whenever the two approaches are non-equivalent, the choice depends on the specific characteristics of the system under study and the nature of the constraints involved. For example, in the context of choosing a null model to be compared with real data for, e.g., pattern detection, it has been argued that, when non-equivalence holds, the ensemble choice should be based on a theoretical guiding principle [11, 9, 12, 13, 14]. In this case, if the data are expected to be noisy, meaning that the measured values of the constraints generally do not match the true (unobservable) ones, one should prefer the canonical version of the null model because it allows for fluctuations in the constraint quantities, while the microcanonical one would paradoxically assign zero probability to the true configuration of the system. Conversely, if, by hypothesis, the measured values of the constraints are not affected by error or noise, the microcanonical model should be preferred.
@@
Information theory provides a powerful framework for model selection by offering criteria that capture the trade-off between a model ‘goodness of fit,’ quantified by its log-likelihood, and a model ‘complexity.’ The two historically most popular model selection criteria are the Akaike information criterion (AIC) [15] and the Bayesian information criterion (BIC) [16], according to which one should pick the model minimizing
@@
respectively, where $\mathcal{L}$ is the model likelihood, $k$ is the number of its parameters and $V$ is the sample size. Canonical and microcanonical models defined by the same set of constraints share the same number of parameters. Therefore, according to both AIC and BIC, their complexity is the same, and a comparison based on these criteria would reduce to a mere comparison between their log-likelihood functions. This would naively lead to the conclusion that the microcanonical model is always the best-scoring one since it has the highest likelihood. However, neither AIC nor BIC can be computed for microcanonical models since they are derived under regularity assumptions that are not even remotely met by the latter. Moreover, both criteria are asymptotic results derived under the assumption of a finite number of parameters as the system size approaches infinity [17]. This assumption is not met by models showing non-equivalence, whose number of parameters scales with the size of the system.
@@
To overcome these issues, here we consider the Minimum Description Length (MDL) principle [18, 19, 20], a family of information theoretical approaches to model selection seeking the model that provides the shortest description of data. Specifically, we focus on the approach based upon the Normalized Maximum Likelihood (NML), or Shtarkov, distribution [21], our aim being that of investigating the interplay between the difference of log-likelihoods and complexities in light of ensemble non-equivalence. Indeed, while the impact of non-equivalence on the difference between log-likelihoods is quite well understood, its effects on the difference between complexities remain an unexplored topic within this framework. To encompass a broad spectrum of applications, our study focuses on maximum-entropy models suitable for studying systems represented by binary rectangular matrices (e.g., bipartite networks, time series). More precisely, we aim at determining i) whether the influence of non-equivalence persists even when the complexity term is accounted for and ii) if non-equivalent canonical and microcanonical models should, in fact, be considered as different from a model selection perspective. We conclude that the answer to both questions is yes, raising the question of whether non-equivalence can be defined on a model selection level. Moreover, as a consequence of our results, we show that, in a Bayesian context, different priors can result in vastly different description lengths when an extensive number of constraints is enforced.
@@
The rest of the paper is organized as follows. Section II introduces MDL for discrete data and the Normalized Maximum Likelihood (NML) approach to define description lengths. In Section III, we provide a general expression for the difference between the description length of canonical and microcanonical models as a function of the Kullback-Leibler divergence between the two distributions. In Section IV, we explicitly derive and compare the description length of both the canonical and the microcanonical formulations of models defined by one global constraint (i.e., the sum of the entries of a matrix) as well as one-sided local constraints (i.e., the row-specific sums of a matrix). While the first constraint leads to ensemble equivalence, the second set of constraints leads to ensemble non-equivalence. Finally, Section V reviews the relationship between the NML-based and Bayesian approaches when non-equivalence holds.
@@ II The Minimum Description Length principle
The MDL principle embodies the idea that the best model to describe a given dataset is the one providing the shortest description. We now introduce the basic formalism by focusing on the discrete data $\bm{x}\in\mathcal{X}$, where $\mathcal{X}$ represents the space of all possible samples of fixed size $V$. A parametric statistical model $\mathcal{M}$ consists of a family of probability distributions sharing the same functional form, i.e.
=$\mathcal{M}=\{P_{\mathcal{M}}(\bm{x};\bm{\theta})\}_{\bm{\theta}\in\Theta}$ (3)
@@
In order to apply the MDL principle, one needs to compute the description length of the data $\bm{x}$ provided by a model $\mathcal{M}$. By Kraft’s inequality, the optimal number of natural digits (nats) needed to describe the outcome $\bm{x}$ of a probability distribution $P(\bm{x})$ is given by $-\ln P(\bm{x})$. Since, however, $\mathcal{M}$ encompasses many distributions corresponding to different parameter values, a natural choice is that of considering the description length provided by the distribution minimizing the quantity $-\ln P_{\mathcal{M}}(\bm{x};\bm{\theta})$, reading
@@ II.1 The Normalized Maximum Likelihood universal distribution
The NML distribution corresponds to the distribution attaining the minimum point-wise regret in the worst-case scenario [18], i.e.
@@
In words, it is the distribution that a priori requires the minimum number of nats when compared to the shortest description length a posteriori, that is $-\ln P_{\mathcal{M}}(\bm{x};\hat{\bm{\theta}}(\bm{x}))$. The minimum is found in the worst-case scenario, i.e.,for the data $\bm{x}$ for which this difference is maximal. The unique solution to this minimax problem reads
@@
Similarly to other criteria rooted in information theory, this description length consists of two terms: the log-likelihood term, evaluated on the data and favouring models with a high goodness-of-fit, and the complexity term, depending only on the chosen model and penalizing the ones that are too complex.
@@ III Description length of canonical and microcanonical models
and denote by $\mathcal{G}$ the set of all such matrices that, when endowed with a probability distribution $P(\mathbf{G})$, constitutes a proper ensemble of matrices. The matrix representation is often used to describe systems composed of $n$ elements to be modelled (corresponding to the number of rows), each characterized by $m$ state variables or degrees of freedom (corresponding to the number of columns). The size of the data set is fixed and supposed to represent the number of independent entries of the matrix, i.e.,$nm$.
@@
where $\mathcal{L}(\mathbf{c})$ denotes the value of the maximum likelihood in correspondence of a sufficient statistics reading $\mathbf{c}$ and $\Omega(\mathbf{c})$ denotes the number of configurations with $\mathbf{c}(\mathbf{G})=\mathbf{c}$, i.e.
@@ III.1 Description length of microcanonical models
where $\mathbf{c}$ is the vector of parameters whose values lie in the discrete set $\mathcal{C}$ for microcanonical models, in fact, the vector of parameters corresponds to the sufficient statistics itself - and $P_{\textrm{mic}}(\mathbf{G};\mathbf{c})$ is a uniform probability whose mass differs from zero only in correspondence of those configurations that verify the condition $\mathbf{c}(\mathbf{G})=\mathbf{c}$.
@@ III.2 Description length of canonical models
with $Z(\bm{\theta})=\sum_{\mathbf{G}\in\mathcal{G}}e^{-\bm{\theta}\cdot\mathbf{c}(\mathbf{G})}$ being the partition function. If the system under study is a network, the maximum-entropy canonical models are also known as Exponential Random Graph Models (ERGMs). In this case, the vector of parameters $\bm{\theta}$ corresponds to the vector of Lagrange multipliers needed to carry out a constrained entropy maximization. The ML estimators $\bm{\hat{\theta}}(\mathbf{G})$ are found by solving the system of equations
@@
$\text{COMP}_{\textrm{can}}^{\text{NML}}$ $=\ln\sum_{\mathbf{c}\in\mathcal{C}}\Omega(\mathbf{c})\cdot\frac{1}{\Omega(\mathbf{c})}\cdot e^{-D_{\text{KL}}(\mathbf{c})}=\ln\sum_{\mathbf{c}\in\mathcal{C}}e^{-D_{\text{KL}}(\mathbf{c})}.$ (25)
As a consequence, the description length of a canonical model can be expressed as
@@ III.3 Comparing canonical and microcanonical models
$Q_{\text{can}}(\mathbf{c};\hat{\bm{\theta}}(\mathbf{c}))$ $=\Omega(\mathbf{c})P_{\text{can}}(\mathbf{c};\hat{\bm{\theta}}(\bm{c}))=\frac{\mathcal{L}_{\text{can}}(\mathbf{c})}{\mathcal{L}_{\text{mic}}(\mathbf{c})}=e^{-D_{\text{KL}}(\mathbf{c})}$ (31)
@@
$\Delta\text{DL}(\mathbf{G}^{*})=-\ln Q_{\text{can}}(\mathbf{c}^{*};\bm{\theta}^{*})+\ln\frac{\sum_{\mathbf{c}\in\mathcal{C}}Q_{\text{can}}(\mathbf{c};\hat{\bm{\theta}}(\mathbf{c}))}{\mathcal{N}_{\mathcal{C}}};$ (32)
@@
$\Delta\text{DL}(\mathbf{G}^{*})=\ln\frac{\bar{Q}_{\text{can}}}{Q_{\text{can}}^{*}}$ (34)
@@
where $o(x)$ stands for a quantity that goes to 00 when divided by $x$ as $n\to+\infty$.
@@
$D_{\text{KL}}(\mathbf{c})=o(n)\quad\forall\mathbf{c}\in\mathcal{C}\quad\Rightarrow\quad\Delta\text{DL}(\mathbf{c})=o(n)\quad\forall\mathbf{c}\in\mathcal{C}$ (37)
@@ IV.1 Enforcing one global constraint
i.e.,the distribution of a collection of $nm$ i.i.d. Bernoulli variables, indicating that each matrix element is either $1$, with probability $p$, or 00, with probability $1-p$. The ML estimator of $p$ is given by the matrix density $\hat{p}(\mathbf{G})=l(\mathbf{G})/nm$ and the maximum log-likelihood, obtained by inserting $\hat{p}$ into (48), reads
@@
$\text{DL}_{\text{can}}^{\text{NML}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)+\ln\left[\frac{e^{nm}\Gamma(nm,nm)}{nm^{nm-1}}+1\right].$ (52)
@@
So far, all the reported results are exact. Nevertheless, as we are interested in the order of $\Delta\text{DL}$ when $n\to\infty$, we consider the following asymptotic results, derived from the exact ones (see section S1.1 of the supplementary material):
@@
The expressions above provide two significant insights. First, the order of the difference between log-likelihood functions is $o(n)$, regardless of the growth rate of $m$, a result implying that condition (36) is met and ensemble equivalence holds. Second, the leading terms of expressions (53) and (54) are identical, thus cancelling out when compared. Consequently, the (asymptotic) difference between description lengths depends on the size of the system only through the matrix density $p^{*}$, assumed to be finite in the dense regime. Thus, such a difference is finite in the same regime, in agreement with (37).

### v4

@@ Abstract
The (non-)equivalence of canonical and microcanonical ensembles is a fundamental question in statistical physics, concerning whether the use of soft and hard constraints in the maximum-entropy construction leads to the same description of a system. Despite the fact that maximum-entropy models are also commonly used in statistical inference, pattern detection, and hypothesis testing, a complete understanding of the effects of ensemble non-equivalence on statistical modeling is still missing. Here, we study this problem from a rigorous model selection perspective by comparing canonical and microcanonical models via the Minimum Description Length (MDL) principle, which yields a trade-off between likelihood, measuring model accuracy, and complexity, measuring model flexibility and its potential to overfit data. We compute the Normalized Maximum Likelihood (NML) of both formulations and find that: (i) microcanonical models always achieve higher likelihood but are always more complex; (ii) the optimal model choice depends on the empirical values of the constraints – the canonical model performs best when its fit to the observed data exceeds its uniform average fit across all realizations; (iii) in the thermodynamic limit, the difference in description length per node vanishes when ensemble equivalence holds but persists otherwise, showing that non-equivalence implies extensive differences between large canonical and microcanonical models. Finally, we compare the NML approach to Bayesian methods, showing that (iv) the choice of priors, practically irrelevant in equivalent models, becomes crucial when an extensive number of constraints is enforced, possibly leading to very different outcomes.
@@ I Introduction
The concept of ensemble equivalence, introduced by Gibbs [6], plays an important role in statistical mechanics. Traditional studies have shown that canonical and microcanonical ensembles tend to become equivalent in the thermodynamic limit for systems with short-range interactions. Violations of this asymptotic equivalence, known as ensemble non-equivalence, have historically been linked to systems with long-range interactions or other forms of non-additivity, such as mean-field spin models or self-gravitating systems [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
More recently, it has become clear that ensemble non-equivalence can also arise due to an additional mechanism: the presence of an extensive number of constraints [19, 20, 21, 22]. This has been demonstrated, for example, in network models with local constraints, such as the Configuration Model. In these cases – unlike the classical examples from statistical physics – non-equivalence is not necessarily associated with long-range interactions or phase transitions, and is not confined to specific parameter regimes, but is present across the entire parameter space. Moreover, this phenomenon can have practical implications for the expected macroscopic properties of the system under study (see [23] for an example concerning bipartite networks).
Whenever the two approaches are equivalent, they identify the same set of typical configurations. In these cases, preferring one formulation over the other becomes a matter of convenience (e.g., computational efficiency). Whenever the two approaches are non-equivalent, the choice depends on the specific characteristics of the system under study and the nature of the constraints involved. For example, in the context of choosing a null model to be compared with real data, for e.g., pattern detection, it has been argued that, when non-equivalence holds, the ensemble choice should be based on a theoretical guiding principle [24, 21, 25, 22, 26]. In this case, if the data are expected to be noisy, meaning that the measured values of the constraints generally do not match the true (unobservable) ones, one should prefer the canonical version of the null model because it allows for fluctuations in the constraint quantities, while the microcanonical one would paradoxically assign zero probability to the true configuration of the system. Conversely, if, by hypothesis, the measured values of the constraints are not affected by error or noise, the microcanonical model should be preferred.
@@
Information theory provides a powerful framework for model selection by offering criteria that capture the trade-off between a model’s “goodness of fit,” quantified by its log-likelihood, and its “complexity.” In general, model complexity measures the flexibility of a model in fitting different datasets. In other words, it measures how easily the model can overfit the data. A common way to quantify model complexity is by counting the number $k$ of the model’s free parameters, which, in the case of maximum-entropy models, corresponds to the number of constraints. This is the case of the two historically most popular model selection criteria: the Akaike Information Criterion (AIC) [27] and the Bayesian Information Criterion (BIC) [28], according to which one should pick the model minimizing
@@
respectively, where $\mathcal{L}$ is the model likelihood and $V$ is the sample size. Canonical and microcanonical models defined by the same set of constraints share the same number of parameters. Therefore, according to both AIC and BIC, their complexity is the same, and a comparison based on these criteria would reduce to a mere comparison between their log-likelihood functions. This would naively lead to the conclusion that the microcanonical model is always the best-scoring one since it has the highest likelihood. However, neither AIC nor BIC can be computed for microcanonical models since they are derived under regularity assumptions that are not even remotely met by the latter. Moreover, both criteria are asymptotic results derived under the assumption of a finite number of parameters as the system size approaches infinity [29]. This assumption is not met by models whose number of parameters scales with the size of the system.
To overcome these issues, here we consider the Minimum Description Length (MDL) principle [30, 31, 32], a family of information-theoretic approaches to model selection seeking the model that provides the shortest description of data. Specifically, we focus on the approach based on the Normalized Maximum Likelihood (NML), or Shtarkov, distribution [33], with the aim of investigating the trade-off between the difference of log-likelihoods and that of complexities in light of ensemble non-equivalence.
A particularly useful definition of non-equivalence, naturally connected to information theory, is the measure-level one, which involves the Kullback-Leibler (KL) divergence between the microcanonical and canonical probability distributions. Under mild conditions, this approach is equivalent to other definitions commonly used in statistical physics [18]. According to this definition, (non-)equivalence is characterized by a (non-)vanishing relative entropy density as the system size approaches infinity. Interestingly, the KL divergence between canonical and microcanonical distributions coincides with the difference between the corresponding log-likelihood functions [19, 20]. Thus, measure-level (non-)equivalence can be inspected by simply considering the asymptotic behavior of this log-likelihood difference [19], providing a natural connection to statistical inference.
@@
While the impact of non-equivalence on the difference between log-likelihoods is relatively well understood [19, 21], its effects on the difference between complexities remain largely unexplored. To encompass a broad spectrum of applications, our study focuses on maximum-entropy models suitable for studying systems represented by binary rectangular matrices (e.g., bipartite networks, time series). More precisely, we aim at determining i) whether the influence of non-equivalence persists even when the complexity term is accounted for, and ii) whether non-equivalent canonical and microcanonical models should, in fact, be regarded as statistically distinct models. We conclude that the answer to both questions is yes, raising the question of whether non-equivalence can be defined at the model selection level. Furthermore, we show that in the presence of extensively many constraints, the choice of prior in a Bayesian framework can critically affect the resulting description length, underscoring the importance of this choice in non-equivalent settings.
@@
The rest of the paper is organized as follows. Section II introduces MDL for discrete data and the Normalized Maximum Likelihood (NML) approach to define description lengths. In Section III, we provide a general expression for the difference between the description lengths of canonical and microcanonical models as a function of the Kullback-Leibler divergence between the two distributions. In Section IV, we explicitly derive and compare the description lengths of both the canonical and the microcanonical formulations of models defined by one global constraint (i.e., the sum of the entries of a matrix) as well as one-sided local constraints (i.e., the row-specific sums of a matrix). While the first constraint leads to ensemble equivalence, the second set of constraints leads to ensemble non-equivalence. Finally, Section V reviews the relationship between the NML-based and Bayesian approaches when non-equivalence holds.
@@ II The Minimum Description Length principle
The MDL principle embodies the idea that the best model to describe a given dataset is the one providing the shortest description. We now introduce the basic formalism by focusing on the discrete data $\bm{x}\in\mathcal{X}$, where $\mathcal{X}$ represents the space of all possible samples of fixed size $V$. A parametric statistical model $\mathcal{M}$ consists of a family of probability distributions sharing the same functional form, i.e.,
$\mathcal{M}=\{P_{\mathcal{M}}(\bm{x};\bm{\theta})\}_{\bm{\theta}\in\Theta}$ (3)
@@
In order to apply the MDL principle, one needs to compute the description length of the data $\bm{x}$ provided by a model $\mathcal{M}$. By Kraft’s inequality [30, 34], the optimal number of natural digits (nats) needed to describe the outcome $\bm{x}$ of a probability distribution $P(\bm{x})$ is given by $\ln P(\bm{x})$. Since, however, $\mathcal{M}$ encompasses many distributions corresponding to different parameter values, a natural choice is that of considering the description length provided by the distribution minimizing the quantity $\ln P_{\mathcal{M}}(\bm{x};\bm{\theta})$, reading
@@ II.1 The Normalized Maximum Likelihood universal distribution
The NML distribution corresponds to the distribution attaining the minimum point-wise regret in the worst-case scenario [30], i.e.,
@@
In words, it is the distribution that a priori requires the minimum number of nats when compared to the shortest description length a posteriori, that is $\ln P_{\mathcal{M}}(\bm{x};\hat{\bm{\theta}}(\bm{x}))$. The minimum is found in the worst-case scenario, i.e., for the data $\bm{x}$ for which this difference is maximal. The unique solution to this minimax problem reads
@@
Similarly to other criteria rooted in information theory, this description length consists of two terms: the log-likelihood term, evaluated on the data and favoring models with a high goodness-of-fit, and the complexity term, depending only on the chosen model and penalizing the ones that are too complex.
@@ III Description length of canonical and microcanonical models
and denote by $\mathcal{G}$ the set of all such matrices that, when endowed with a probability distribution $P(\mathbf{G})$, constitute a proper ensemble of matrices. The matrix representation is often used to describe systems composed of $n$ elements to be modeled (corresponding to the number of rows), each characterized by $m$ state variables or degrees of freedom (corresponding to the number of columns). The size of the data set is fixed and supposed to represent the number of independent entries of the matrix, i.e., $nm$.
@@
where $\mathcal{L}(\mathbf{c})$ denotes the value of the maximum likelihood in correspondence with a sufficient statistic reading $\mathbf{c}$ and $\Omega(\mathbf{c})$ denotes the number of configurations with $\mathbf{c}(\mathbf{G})=\mathbf{c}$, i.e.,
@@ III.1 Description length of microcanonical models
where $\mathbf{c}$ is the vector of parameters whose values lie in the discrete set $\mathcal{C}$ - for microcanonical models, in fact, the vector of parameters corresponds to the sufficient statistics itself - and $P_{\textrm{mic}}(\mathbf{G};\mathbf{c})$ is a uniform probability whose mass differs from zero only in correspondence of those configurations that verify the condition $\mathbf{c}(\mathbf{G})=\mathbf{c}$. The derivation of this probability functional form as the solution of entropy maximization under hard constraints is given in more detail in Section A of the Appendix.
@@ III.2 Description length of canonical models
with $Z(\bm{\theta})=\sum_{\mathbf{G}\in\mathcal{G}}e^{-\bm{\theta}\cdot\mathbf{c}(\mathbf{G})}$ being the partition function. In the context of network theory, the maximum-entropy canonical models are also known as Exponential Random Graph Models [36, 37, 2]. In this case, the vector of parameters $\bm{\theta}$ corresponds to the vector of Lagrange multipliers needed to carry out a constrained entropy maximization. The ML estimators $\bm{\hat{\theta}}(\mathbf{G})$ are found by solving the system of equations
@@
$\text{COMP}_{\textrm{can}}^{\text{NML}}$ $=\ln\sum_{\mathbf{c}\in\mathcal{C}}\Omega(\mathbf{c})\cdot\frac{1}{\Omega(\mathbf{c})}\cdot e^{-D_{\text{KL}}(\mathbf{c})}$
$=\ln\sum_{\mathbf{c}\in\mathcal{C}}e^{-D_{\text{KL}}(\mathbf{c})}.$ (25)
Consequently, the description length of a canonical model can be expressed as
@@ III.3 Comparing canonical and microcanonical models
$Q_{\text{can}}(\mathbf{c};\hat{\bm{\theta}}(\mathbf{c}))$ $=\Omega(\mathbf{c})P_{\text{can}}(\mathbf{c};\hat{\bm{\theta}}(\bm{c}))$
$=\frac{\mathcal{L}_{\text{can}}(\mathbf{c})}{\mathcal{L}_{\text{mic}}(\mathbf{c})}=e^{-D_{\text{KL}}(\mathbf{c})}$ (31)
@@
$\Delta\text{DL}(\mathbf{G}^{*})=-\ln Q_{\text{can}}(\mathbf{c}^{*};\bm{\theta}^{*})\\ +\ln\frac{\sum_{\mathbf{c}\in\mathcal{C}}Q_{\text{can}}(\mathbf{c};\hat{\bm{\theta}}(\mathbf{c}))}{\mathcal{N}_{\mathcal{C}}};$ (32)
@@
$\Delta\text{DL}(\mathbf{G}^{*})=\ln\frac{\bar{Q}_{\text{can}}}{Q_{\text{can}}^{*}},$ (34)
@@
where $o(x)$ stands for a quantity that goes to 0 when divided by $x$ as $n\to+\infty$.
@@
$D_{\text{KL}}(\mathbf{c})=o(n)\quad\forall\mathbf{c}\in\mathcal{C}\\ \quad\Rightarrow\quad\Delta\text{DL}(\mathbf{c})=o(n)\quad\forall\mathbf{c}\in\mathcal{C}$ (37)
@@ IV.1 Enforcing one global constraint
i.e., the distribution of a collection of $nm$ i.i.d. Bernoulli variables, indicating that each matrix element is either $1$, with probability $p$, or 0, with probability 1p1-p. The ML estimator of $p$ is given by the matrix density $\hat{p}(\mathbf{G})=l(\mathbf{G})/nm$ and the maximum log-likelihood, obtained by inserting $\hat{p}$ into (48), reads
@@
$\text{DL}_{\text{can}}^{\text{NML}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)\\ +\ln\left[\frac{e^{nm}\Gamma(nm,nm)}{nm^{nm-1}}+1\right].$ (52)
@@
So far, all the reported results are exact. Nevertheless, as we are interested in the order of $\Delta\text{DL}$ when $n\to\infty$, we consider the following asymptotic results, derived from the exact ones (see section B.1 of the Appendix):
@@
The expressions above provide two significant insights. First, the order of the difference between log-likelihood functions is $o(n)$, regardless of the growth rate of $m$, a result implying that condition (36) is met and ensemble equivalence holds. Second, the leading terms of expressions (53) and (54) are identical, thus canceling out when compared. Consequently, the (asymptotic) difference between description lengths depends on the size of the system only through the matrix density $p^{*}$, assumed to be finite in the dense regime. Thus, such a difference is finite in the same regime, in agreement with (37).